# From Chat Logs to Collective Insights: Aggregative Question Answering

### v1

@@ Abstract
Conversational agents powered by large language models (LLMs) are rapidly becoming integral to our daily interactions, generating unprecedented amounts of conversational data. Such datasets offer a powerful lens into societal interests, trending topics, and collective concerns. Yet, existing approaches typically treat these interactions as independent and miss critical insights that could emerge from aggregating and reasoning across large-scale conversation logs. In this paper, we introduce Aggregative Question Answering, a novel task requiring models to reason explicitly over thousands of user-chatbot interactions to answer aggregative queries, such as identifying emerging concerns among specific demographics. To enable research in this direction, we construct a benchmark, WildChat-AQA, comprising 6,027 aggregative questions derived from 182,330 real-world chatbot conversations. Experiments show that existing methods either struggle to reason effectively or incur prohibitive computational costs, underscoring the need for new approaches capable of extracting collective insights from large-scale conversational data.
@@ 1 Introduction
Rapid adoption of conversation agents powered by large language models (LLMs) is transforming human-computer interactions, integrating deeply into society, and generating unprecedented volumes of conversational data (Backlinko Team, 2025; Vynck, 2023). Platforms using LLM-based chatbots now routinely handle millions of interactions daily, producing rich datasets that capture real-time dialogues reflecting genuine user interests, emerging societal trends, and collective concerns (Zhao et al., 2024b; Zheng et al., 2024). Such conversational data offer immense potential for deriving insights at scale, revealing patterns in societal dynamics, shifts in public sentiment, and demographic-specific concerns.
@@
To address this gap, we introduce a new task, Aggregative Question Answering, which requires reasoning across large-scale collections of user-chatbot interactions to extract aggregative insights. Unlike traditional summarization, which condenses information from one or a few documents into static summaries, Aggregative Question Answering generates dynamic answers that explicitly depend on the specific aggregative query posed. The task requires holistic reasoning over thousands of conversations to answer questions such as identifying trending topics within specific timeframes (What topics trended last week?”), emerging concerns among particular demographics (What topics are Californians concerned about before an election?”), or tracking changes in societal sentiment (“How have users’ attitudes toward artificial intelligence evolved this month?”). The core challenge thus lies not in summarizing individual conversations, but rather in global-scale reasoning conditioned on the query. Figure 1 highlights the high-level distinctions between traditional summarization, aggregative question answering, and querying predefined databases.
@@
To facilitate research into Aggregative Question Answering, we introduce a benchmark WildChat-AQA, constructed from the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat captures not only conversation transcripts but also metadata such as temporal, geographical, and user-specific information. WildChat-AQA formulates aggregational queries about both explicit and implicit attributes of conversations—including topics, keywords, geographical locations, and time—in a multi-choice format. A concrete example of the data creation process is shown in Figure 2. The benchmark comprises 6,027 aggregative questions derived from 182,330 real-world user-chatbot conversations, reflecting genuine user interests and societal trends, thus providing a resource for evaluating models’ ability to reason holistically at scale.
@@
We evaluate current methods, including both non-reasoning and reasoning models, adapted to this task via fine-tuning, retrieval-augmented generation (RAG), and a customized retrieval approach developed specifically for aggregative reasoning: PROBE (Probing Retrieval Of Broad Evidence). Experimental results reveal substantial limitations in existing methods: current systems either struggle to reason effectively at scale or incur prohibitive computational costs. Even under idealized conditions—when the exact oracle contexts relevant to a query are provided—significant room for improvement remains; under realistic conditions, the performance further deteriorates.
@@
Our findings emphasize the critical need for more scalable and effective methods capable of extracting collective insights from large-scale conversational datasets. While Aggregative Question Answering opens promising avenues for impactful real-world analytics, we acknowledge potential societal impacts, particularly when insights relate to sensitive topics like elections, public opinion, or public health—areas that could potentially be susceptible to manipulation. Nevertheless, we strongly believe that transparent, open academic research fosters responsible development and deployment of such powerful technologies. By introducing Aggregative Question Answering as a new task, we aim to spur future methods that fully harness the potential of large-scale conversational data, ultimately enabling deeper societal understanding and more responsive, informed applications of LLMs.
@@
Our benchmark, code, and dataset are publicly available at https://github.com/yuntian-group/wildchat_aqa, and we also provide a user-friendly benchmark visualization tool at http://65.108.32.135:3000/dataview.
2 WildChat-AQA
@@
To support research on Aggregative Question Answering, we constructed the WildChat-AQA benchmark based on the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat provides real-world conversations between users and chatbots, along with basic metadata such as timestamps and user locations. In this work, we extended these attributes by introducing additional attributes such as topics and keywords inferred from the conversation text using LLMs. These inferred attributes serve as the ground-truth annotations for constructing our benchmark. At evaluation time, models must infer them from conversations to answer aggregative questions. Table 1 summarizes the attributes and indicates which ones require inference and which ones are directly available.
| Name | Multi-Val | Inferred | Example |
| Location | No | No | United States, Canada |
| User Name | No | No | lostclasp37, toughcue8 |
| Time | No | No | 4/26/2023, 1:47:24 PM |
| Language | No | No | English, Russian |
| Topic | Yes | Yes | Software, Programming and Computer Science |
| Subtopic | Yes | Yes | Mobile Development, AI and ML |
@@ 2.1 Dataset Construction
The construction of WildChat-AQA involved four main steps, as illustrated in Figure 2:
@@
We began by performing minHash-based deduplication (Hugging Face, 2023) to remove highly similar conversations, ensuring diversity. We also filtered conversations exceeding 4,096 tokens to maintain manageable context lengths. Additionally, we retained only active users (those with at least 10 interactions) to ensure sufficient user-specific data. We also generated user IDs from IP addresses and headers.
@@
To support meaningful aggregative queries, we prompted GPT-4o to summarize each conversation and extract relevant keywords. Using these summaries, we recursively applied TnT-LLM (Wan et al., 2024) to infer hierarchical topics at two levels: coarse-grained topics and fine-grained subtopics. Detailed prompts and examples can be found in Appendix E.
Step 3: Keywords Categorization
Certain subtopics, such as “Programming” and “Fan-fiction and Crossover,” contained many conversations. To support finer-grained aggregative queries, we further categorized keywords inferred from conversations into higher-level categories using LLMs so that we can derive aggregative information. For example, different Pokémon-related keywords (versions, characters, trademarks) were grouped into a single category “Pokémon”. Full details of this procedure are also available in Appendix E.
@@
Finally, we generated aggregative questions using combinations of attributes stored in our constructed database. We systematically sampled attribute combinations (from zero to three attributes as conditions) to query our database for a remaining target attribute. To ensure questions were both diverse and answerable, attribute combinations were carefully sampled. These structured queries were then translated into natural language using GPT-4.1.
@@ 2.2 Dataset Statistics
The resulting WildChat-AQA benchmark contains 182,330 user-chatbot conversations and 6,027 aggregative questions. These conversations cover 28 high-level topics, 455 fine-grained subtopics, and 14,482 keyword categories. Detailed statistics of questions organized by different attribute conditions and target attributes are shown in Table 8 in Appendix B. Unlike typical question-answering tasks, which derive answers from one or a few documents, WildChat-AQA requires models to reason over contexts whose total token counts range widely from to tokens. Figure 4 illustrates the distribution of context token counts. Full data statistics are provided in Appendix B.
@@ 2.3 Evaluation Protocol
We frame the evaluation of aggregative question answering as a ranking problem. During training, the model or system under evaluation is provided access to the entire WildChat-AQA dataset. At test time, the model is given an aggregative question along with 10 candidate answers. Its task is to rank these candidates according to their relevance to the question. We use standard ranking metrics NDCG@1, NDCG@3, NDCG@5, and NDCG@10 to measure performance.
@@ 2.4 Human Evaluation
To evaluate the quality of our inferred attributes, we conducted a human evaluation measuring both inter-annotator agreement (human-human) and human-model agreement using Cohen’s . Specifically, we randomly sampled 100 examples each for level-1 (topic) and level-2 (subtopic) taxonomy labeling. Due to the multi-label nature of these tasks, we computed per-label agreement by treating each possible category as an independent binary labeling task. For subtopic evaluation, we additionally reported macro-average agreement scores aggregated across all topics to provide a comprehensive view of annotation reliability.
@@ 3 Probing Retrieval Of Broad Evidence
Traditional retrieval methods, including those used in retrieval-augmented generation (RAG), typically aim to identify a small set of highly specific, relevant documents. However, for Aggregative Question Answering, it is essential to identify a broader range of documents that collectively support reasoning about high-level aggregational insights. To address this unique requirement, we introduce a customized retrieval method, Probing Retrieval Of Broad Evidence (PROBE). PROBE operates in two main steps:
@@
Given a question , we first prompt a large language model to generate a comprehensive set of short, diverse queries that may help retrieve a broad range of relevant documents. Specifically, the LLM generates a set of queries related to the question. Additionally, the model generates strict filtering conditions to exclude documents clearly unrelated to the question. Formally, this process is defined as:
@@
Next, each generated query (alongside the filtering conditions ) is used individually to retrieve relevant documents. This results in separate retrieval runs. We then aggregate these results by merging the retrieved document lists according to their retrieval relevance scores. If a document appears multiple times across different queries, we use max pooling to assign it the highest relevance score it received from any query. Finally, we select the top documents from this aggregated list as evidence.
The resulting set of retrieved documents serves as supporting evidence for the model to perform aggregational reasoning and answer the question. An overview of the full PROBE retrieval pipeline is illustrated in Figure 3.
@@ 4 Experiments
We evaluate WildChat-AQA using various models, retrieval approaches, and conversation formats. We also explore model behaviors by performing ablation studies.
@@ 4.2 Experimental Setups
No Context
The model directly answers questions without external inputs, relying solely on internal knowledge. This approach establishes baseline performance using only pre-existing knowledge. Due to resource constraints, we only evaluated this baseline using the strongest reasoning model (o4-mini).
@@
We finetune pretrained models on the entire WildChat-AQA dataset.
@@
We use our proposed retrieval method, PROBE. Query generation uses GPT-4.1-mini, and retrieval relies on embeddings from OpenAI’s text-embedding-3-large model.
@@
Raw conversations are detailed but noisy (average 1,143.4 tokens each), whereas summarized conversations are more concise (average 21.5 tokens). Thus, we experimented with both raw and summarized conversation inputs to investigate their effectiveness for aggregative question answering. Implementation details for experiments are provided in Appendix D.
@@ 4.4 Main Results
| Model Name | Approach | Type | NDCG@1 | NDCG@3 | NDCG@5 | NDCG@10 | # Input Token (Million) |
| Random | None | / | 0.2501 | 0.3516 | 0.4368 | 0.6211 | - |
| o4-mini | None | / | 0.3063 | 0.4017 | 0.4805 | 0.6488 | 0.87 |
| Qwen3 8B | Finetune | Raw | 0.2694 | 0.3739 | 0.4589 | 0.6346 | 1.74 |
| Summary | 0.2984 | 0.3966 | 0.4807 | 0.6480 | 1.74 | ||
| Gemma3 4B | RAG | Raw | 0.3291 | 0.4356 | 0.5159 | 0.6688 | 73.48 |
| Summary | 0.3740 | 0.4895 | 0.5627 | 0.6991 | 174.62 | ||
| PROBE | Raw | 0.4766 | 0.5891 | 0.6478 | 0.7620 | 38.44 | |
| Summary | 0.5430 | 0.6513 | 0.6994 | 0.7969 | 17.35 | ||
| Qwen3 8B Think | RAG | Raw | 0.4168 | 0.5090 | 0.5779 | 0.7123 | 362.16 |
| Summary | 0.5273 | 0.6110 | 0.6646 | 0.7717 | 176.88 | ||
| PROBE | Raw | 0.6545 | 0.7305 | 0.7728 | 0.8483 | 315.52 | |
| Summary | 0.6944 | 0.7638 | 0.8005 | 0.8660 | 123.04 | ||
| Qwen3 32B Think | RAG | Raw | 0.4052 | 0.5020 | 0.5705 | 0.7081 | 182.90 |
| Summary | 0.5496 | 0.6321 | 0.6847 | 0.7850 | 176.88 | ||
| PROBE | Raw | 0.6525 | 0.7347 | 0.7759 | 0.8501 | 315.52 | |
| Summary | 0.7056 | 0.7753 | 0.8114 | 0.8725 | 123.04 | ||
| GPT-4.1 mini | RAG | Raw | 0.4494 | 0.5387 | 0.6035 | 0.7299 | 344.37 |
| Summary | 0.5782 | 0.6620 | 0.7104 | 0.8019 | 154.31 | ||
| PROBE | Raw | 0.6806 | 0.7536 | 0.7936 | 0.8628 | 298.69 | |
| Summary | 0.7308 | 0.7942 | 0.8282 | 0.8843 | 107.11 | ||
| o4-mini | RAG | Raw | 0.4730 | 0.5510 | 0.6116 | 0.7383 | 344.37 |
| Summary | 0.6122 | 0.6792 | 0.7242 | 0.8140 | 154.31 | ||
| PROBE | Raw | 0.7117 | 0.7747 | 0.8086 | 0.8745 | 298.69 | |
@@
Table 3 presents performance results across different models, retrieval methods, and conversation formats.
@@
Among tested models, o4-mini consistently achieved the highest performance, with a maximum NDCG@1 score of 0.7571. GPT-4.1-mini, while also strong, trailed slightly behind. Among open-source models, Qwen3-32B-think achieves the highest performance (0.7056 NDCG@1).
PROBE improves significantly over standard RAG.
Compared to standard RAG, PROBE consistently showed large performance improvements. On raw data, PROBE improved NDCG@1 scores by 14.8, 23.7, 24.7, 23.1, and 23.8 points for Gemma3-4B, Qwen3-8B-think, Qwen3-32B-think, GPT-4.1-mini, and o4-mini, respectively. A similar trend was observed using summarized conversations.
@@
Models consistently performed better with summarized inputs, showing improved NDCG@1 scores of 4.5 to 14.4 points over raw conversations for standard RAG, and 4.0 to 6.6 points for PROBE. Summaries enable more efficient information retrieval and easier aggregation of insights.
Finetuning doesn’t help.
Finetuning Qwen3-8B did not significantly exceed random-chance performance. This result suggests that current finetuning methods fail to internalize aggregative information effectively from conversational datasets.
High token consumption.
Achieving good performance on this task required models to consume a very large number of input tokens as shown in Table 3. This highlights a significant computational challenge and motivates future research to improve efficiency.
@@
| # Conversation | Approach | Recall | NDCG@5 |
| 5 | RAG | 0.01 | 0.5373 |
| PROBE | 0.07 | 0.6991 | |

### v2

@@ Abstract
Conversational agents powered by large language models (LLMs) are rapidly becoming integral to our daily interactions, generating unprecedented amounts of conversational data. Such datasets offer a powerful lens into societal interests, trending topics, and collective concerns. Yet, existing approaches typically treat these interactions as independent and miss critical insights that could emerge from aggregating and reasoning across large-scale conversation logs. In this paper, we introduce Aggregative Question Answering, a novel task requiring models to reason over thousands of user-chatbot interactions to answer aggregative queries, such as identifying emerging concerns among specific demographics. To enable research in this direction, we constructed WildChat-AQA, a benchmark comprising 6,027 aggregative questions derived from 182,330 real-world chatbot conversations. Experiments show that existing methods either struggle to reason effectively or incur prohibitive computational costs, underscoring the need for new approaches capable of extracting collective insights from large-scale conversational data.
@@ 1 Introduction
Rapid adoption of conversation agents powered by large language models (LLMs) is transforming human-computer interactions, integrating deeply into society, and generating unprecedented volumes of conversational data (Backlinko Team, 2025; Vynck, 2023). Platforms using LLM-based chatbots now routinely handle millions of interactions every day, producing rich datasets that capture real-time dialogues reflecting user interests, emerging societal trends, and collective concerns (Zhao et al., 2024b; Zheng et al., 2024). Such conversational data offer immense potential for deriving insights at scale, revealing patterns in societal dynamics, shifts in public sentiment, and demographic-specific concerns (Valdez et al., 2020).
@@
To address this gap, we introduce a new task, Aggregative Question Answering, which requires reasoning across large-scale collections of user-chatbot interactions to extract aggregative insights. Unlike traditional summarization, which condenses information from one or a few documents into static summaries, Aggregative Question Answering generates dynamic answers that depend on the specific aggregative query posed. The task requires reasoning over thousands of conversations to answer questions such as identifying trending topics within specific timeframes (“What topics trended last week?”), emerging concerns among particular demographics (“What topics are Californians concerned about before an election?”), or tracking changes in societal sentiment (“How have users’ attitudes toward AI evolved this month?”). The core challenge thus lies not in summarizing individual conversations, but rather in global-scale reasoning conditioned on the query. Figure˜1 highlights the high-level distinctions between traditional summarization, querying structured databases, and aggregative question answering.
@@
To facilitate research into Aggregative Question Answering, we introduce WildChat-AQA, a benchmark constructed from the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat captures not only conversation transcripts but also metadata such as temporal, geographical, and user-specific information. WildChat-AQA formulates aggregational queries about both explicit and implicit attributes of conversations, including topics, keywords, geographical locations, and temporal information, in a multi-choice format. A concrete example of the data creation process is shown in Figure˜2. The benchmark includes 6,027 aggregative questions derived from 182,330 real-world user-chatbot conversations, reflecting genuine user interests and societal trends, thus providing a resource for evaluating models’ ability to perform aggregative reasoning at scale.
@@
We evaluated current methods, including both non-reasoning and reasoning models, adapted to this task via fine-tuning, retrieval-augmented generation (RAG), and a customized retrieval approach developed specifically for aggregative reasoning: PROBE (Probing Retrieval Of Broad Evidence). Experimental results show substantial limitations in existing methods: current systems either struggle to reason effectively at scale or incur prohibitive computational costs. Even when whole oracle contexts relevant to a query are provided, there remains significant room for improvement. In more realistic settings with no access to oracle contexts, the performance drops further.
@@
Our findings show that we need more scalable and effective methods capable of extracting collective insights from large-scale conversational datasets. While Aggregative Question Answering opens promising avenues for real-world analytics, we acknowledge potential societal impacts, particularly when insights relate to sensitive topics such as elections, public opinion, or public health. However, we believe that transparent, open academic research fosters responsible development and deployment of such powerful technologies. By introducing Aggregative Question Answering as a new task, we aim to spur future methods that fully harness the potential of large-scale conversational data, ultimately enabling deeper societal understanding and more impactful applications of LLMs.
@@
Our benchmark, code, and dataset are publicly available at https://github.com/yuntian-group/wildchat_aggregative_question_answering, and we also provide a user-friendly benchmark visualization tool at https://aggregativeqa.com/dataview.
2 Aggregative Question Answering
@@
To support research on Aggregative Question Answering, we construct the WildChat-AQA benchmark based on the WildChat dataset (Zhao et al., 2024b; Deng et al., 2024). WildChat provides real-world conversations between users and chatbots, along with basic metadata such as timestamps and user locations. In this work, we extend these attributes by introducing additional attributes, such as topics and keywords inferred from the conversation text using LLMs. These inferred attributes serve as the ground truth annotations for building our benchmark. At evaluation time, models must infer them from conversations to answer aggregative questions. Table˜1 summarizes the attributes, indicating which require inference and which can be obtained directly.
| Name | Multi-Val | Inferred | Examples | ||
| Location | No | No | United States, Canada | ||
| User Name | No | No | lostclasp37, toughcue8 | ||
| Time | No | No | 4/26/2023, 1:47:24 PM | ||
| Language | No | No | English, Russian | ||
| Topic | Yes | Yes |
| Subtopic | Yes | Yes |
@@ 2.1 Dataset Construction
The construction of WildChat-AQA involve four main steps, as illustrated in Figure˜2:
@@
We begin by performing minHash-based deduplication (Hugging Face, 2023) to remove highly similar conversations to ensure diversity. We also filter conversations that exceed 4,096 tokens to maintain manageable context lengths. Additionally, we retain only active users with at least 10 interactions to ensure sufficient user-specific data. We also generate user IDs from IP addresses and headers.
@@
To support meaningful aggregative queries, we prompt GPT-4o to summarize each conversation and extract relevant keywords. Using these summaries, we recursively apply TnT-LLM (Wan et al., 2024) to infer hierarchical topics at two levels: coarse-grained topics and fine-grained subtopics. Detailed prompts and examples can be found in Appendix˜E.
Step 3: Keyword Categorization
Certain subtopics, such as “Programming” and “Fan-fiction and Crossover,” contain many conversations. To support finer-grained aggregative queries, we further categorize keywords inferred from conversations into higher-level categories using LLMs so that we can derive aggregative information. For example, different Pokémon-related keywords (versions, characters, trademarks) are grouped into a single category “Pokémon”. Full details of this procedure are also available in Appendix˜E.
@@
Finally, we generate aggregative questions using combinations of attributes stored in our constructed database. This database is built by compiling all conversations along with their inferred attributes (such as topics and keywords extracted by GPT-4o) and provided metadata attributes (such as timestamps and locations). We then sample attribute combinations from zero to three attributes as conditions and a target attribute to query our database. These structured queries explicitly specify the conditions (attribute-value constraints, e.g., user=abcd, keyword=efgh) and the target attributes to query. The exact combinations are detailed in Table˜8 and Appendix˜B. They serve two purposes: (1) retrieving ground truth answers by converting them directly into database queries executed against our database to retrieve and rank candidate answers, and (2) generating corresponding natural language questions using GPT-4.1. The prompts used are provided in Figure˜24 and Appendix˜E.
@@ 2.2 Dataset Statistics
The final WildChat-AQA benchmark contains 182,330 user-chatbot conversations and 6,027 aggregative questions. These conversations cover 28 high-level topics, 455 fine-grained subtopics, and 14,482 keyword categories. Table˜8 in Appendix˜B provides detailed statistics of the questions organized by different attribute conditions and target attributes. Unlike typical question-answering tasks, which derive answers from one or a few documents, WildChat-AQA requires models to reason over contexts whose total token counts range widely from to tokens. Figure˜4 illustrates the distribution of context token counts. Full data statistics are provided in Appendix˜B.
@@ 2.3 Evaluation Protocol
We frame the evaluation of aggregative question answering as a ranking problem. During training, the model or system is provided access to the entire WildChat-AQA dataset. At test time, the model is given an aggregative question along with 10 candidate answers. Its task is to rank these candidates according to their relevance to the question. We use standard ranking metrics NDCG@1, NDCG@3, NDCG@5, and NDCG@10 to measure performance.
@@ 2.4 Human Evaluation
To evaluate the quality of our inferred attributes, we conduct a human evaluation measuring both inter-annotator agreement (human-human) and human-model agreement using Cohen’s . Specifically, we randomly sample 100 examples each for level-1 (topic) and level-2 (subtopic) taxonomy labeling. Due to the multi-label nature of these tasks, we compute per-label agreement by treating each possible category as an independent binary labeling task. For subtopic evaluation, we additionally report macro-average agreement scores aggregated across all topics to provide a comprehensive view of annotation reliability.
@@ 3 Probing Retrieval Of Broad Evidence
Traditional retrieval methods, including those used in retrieval-augmented generation (RAG) (Lewis et al., 2020), typically aim to identify a small set of highly specific, relevant documents. However, for Aggregative Question Answering, it is essential to identify a broader range of documents that collectively support reasoning about high-level aggregational insights. To address this unique requirement, we introduce a customized retrieval method, Probing Retrieval Of Broad Evidence (PROBE). PROBE operates in two main steps:
@@
Given a question , we first prompt an LLM to generate a comprehensive set of short, diverse queries that may help retrieve a broad range of relevant documents. Specifically, the LLM generates a set of queries related to the question. Additionally, the model generates strict filtering conditions to exclude documents clearly unrelated to the question. Formally, this process is defined as:
@@
Next, each generated query along with the filtering conditions is used individually to retrieve relevant documents. This results in separate retrieval runs. We then aggregate these results by merging the retrieved document lists according to their retrieval relevance scores. If a document appears multiple times across different queries, we use max pooling to assign it the highest relevance score it received from any query. Finally, we select the top documents from this aggregated list as evidence.
The resulting set of retrieved documents serves as supporting evidence for the model to perform aggregational reasoning and answer the question. An overview of the full PROBE retrieval pipeline is in Figure˜3.
@@ 4.2 Experimental Setups
Textual Similarity
We use textual similarity score including BM25 and embedding-based consine similarity using text-embedding-3-large embeddings (denoted as Cosine Sim) to rank the response without context information.
Model With No Context
The model directly answers questions without external inputs, relying solely on internal knowledge. This approach establishes baseline performance using only pre-existing knowledge. Due to resource constraints, we only evaluate this baseline using o4-mini, which is one of the strongest reasoning models.
@@
We finetune pretrained models on the entire WildChat-AQA raw conversations and summaries to test whether fine-tuning on context can bring improvement to the QA tasks.
@@
For our retrieval method, PROBE,, query generation uses GPT-4.1-mini, and the retrieval relies on embeddings from OpenAI’s text-embedding-3-large model.
@@
Raw conversations are detailed but noisy (average 1,143.4 tokens each), whereas summarized conversations are more concise (average 21.5 tokens). Therefore, we experiment with both raw and summarized conversation inputs to investigate their effectiveness for aggregative question answering. Implementation details for experiments are provided in Appendix˜D.
@@ 4.4 Main Results
| Model Name | Context | Type | NDCG@1 | NDCG@3 | NDCG@5 | NDCG@10 |
| Random | - | - | 0.2501 | 0.3516 | 0.4368 | 0.6211 | - | ||
| BM25 | - | - | 0.2320 | 0.3529 | 0.4385 | 0.6208 | - | ||
| Cosine Sim | - | - | 0.2761 | 0.3795 | 0.4638 | 0.6382 | - | ||
| o4-mini | - | - | 0.3063 | 0.4017 | 0.4805 | 0.6488 | 0.87 | ||
| Qwen3 8B | Finetune | Raw | 0.2694 | 0.3739 | 0.4589 | 0.6346 | 1.74 | ||
| Summary | 0.2984 | 0.3966 | 0.4807 | 0.6480 | 1.74 | ||||
| Gemma3 4B | RAG | Raw | 0.3291 | 0.4356 | 0.5159 | 0.6688 | 73.48 | ||
| Summary | 0.3740 | 0.4895 | 0.5627 | 0.6991 | 174.62 | ||||
| PROBE | Raw | 0.4766 | 0.5891 | 0.6478 | 0.7620 | 38.44 | |||
| Summary | 0.5430 | 0.6513 | 0.6994 | 0.7969 | 17.35 | ||||
RAG | Raw | 0.4168 | 0.5090 | 0.5779 | 0.7123 | 362.16 | ||
| Summary | 0.5273 | 0.6110 | 0.6646 | 0.7717 | 176.88 | ||||
| PROBE | Raw | 0.6545 | 0.7305 | 0.7728 | 0.8483 | 315.52 | |||
| Summary | 0.6944 | 0.7638 | 0.8005 | 0.8660 | 123.04 | ||||
RAG | Raw | 0.4052 | 0.5020 | 0.5705 | 0.7081 | 182.90 | ||
| Summary | 0.5496 | 0.6321 | 0.6847 | 0.7850 | 176.88 | ||||
| PROBE | Raw | 0.6525 | 0.7347 | 0.7759 | 0.8501 | 315.52 | |||
| Summary | 0.7056 | 0.7753 | 0.8114 | 0.8725 | 123.04 | ||||
| GPT-4.1 mini | RAG | Raw | 0.4494 | 0.5387 | 0.6035 | 0.7299 | 344.37 | ||
| Summary | 0.5782 | 0.6620 | 0.7104 | 0.8019 | 154.31 | ||||
| PROBE | Raw | 0.6806 | 0.7536 | 0.7936 | 0.8628 | 298.69 | |||
| Summary | 0.7308 | 0.7942 | 0.8282 | 0.8843 | 107.11 | ||||
| o4-mini | RAG | Raw | 0.4730 | 0.5510 | 0.6116 | 0.7383 | 344.37 | ||
| Summary | 0.6122 | 0.6792 | 0.7242 | 0.8140 | 154.31 | ||||
| PROBE | Raw | 0.7117 | 0.7747 | 0.8086 | 0.8745 | 298.69 | |||
@@
Table˜3 presents performance results across different models, retrieval methods, and conversation formats.
Simple textual relevance is ineffective.
We experiment with simple BM25 and embedding-based textual similarity models. We find that textual relevance baselines performed no better than random selection. The embedding-based approach performs slightly better than BM25, improving NDCG@1, 3, 5, and 10 by 4.41, 2.66, 2.53, and 1.74, respectively.
@@
Among tested models, o4-mini consistently achieves the highest performance, with a maximum NDCG@1 score of 0.7571. GPT-4.1-mini, while also strong, trails slightly behind. Among open-source models, Qwen3-32B-think achieves the highest performance. (0.7056 NDCG@1).
PROBE outperforms standard RAG.
Compared to standard RAG, PROBE consistently shows large performance improvements. With raw data, PROBE improves NDCG@1 scores by 14.8, 23.7, 24.7, 23.1, and 23.8 points for Gemma3-4B, Qwen3-8B-think, Qwen3-32B-think, GPT-4.1-mini, and o4-mini, respectively. A similar trend is observed using summarized conversations.
@@
Models consistently perform better with summarized inputs, showing improved NDCG@1 scores of 4.5 to 14.4 points over raw conversations for standard RAG, and 4.0 to 6.6 points for PROBE. Summaries enable more efficient information retrieval and easier aggregation of insights.
Basic finetuning is not effective.
Direct finetuning on Qwen3-8B (raw or summarized conversations without explicit aggregative reasoning steps) does not substantially exceed random-chance performance. This suggests that standard finetuning alone may be insufficient to internalize aggregative information. We caution, however, against generalizing this finding to all finetuning strategies: more sophisticated approaches that explicitly incorporate aggregative reasoning traces during training could yield stronger results, making this an important avenue for future work.
Token consumption is high.
Achieving good performance on this task requires models to consume a very large number of input tokens as shown in Table˜3. This highlights a significant computational challenge and motivates future research to improve efficiency.
@@
| # Conversation | Context | Recall | NDCG@5 | # Input Token (M) |
| 5 | RAG | 0.01 | 0.5373 | 0.33 |
| PROBE | 0.07 | 0.6991 | 0.33 | |